OpenStudy (anonymous):

Write an equation of the line containing the given point and parallel to the given line.

OpenStudy (anonymous):

*8,9); x+7y = 8

OpenStudy (anonymous):

(8,9): x + 7y = 8

OpenStudy (anonymous):

you will want to set it up as y = ax + b. parallel means that they share the same slope [a]. by plugging in 8 and 9 as x and y respectively, you can find a.